# Marginal treatment effects in the absence of instrumental variables

### v1

@@ 3 Identification
Our identification strategy relies on a linearity restriction on the potential outcome equations and a nonlinearity restriction on the propensity score function. The intuition is that the propensity score minus the linear outcome index will provide the excluded variation perturbing treatment status, which plays the role of a continuous IV. Furthermore, to ensure the exogeneity of the nonlinear variation, it is necessary to impose a certain form of independence between the covariates and potential outcome residuals. Specifically, we make the following assumptions to achieve the identification of $\Delta^{\text{MTE}}\left(x,v\right)$. Before proceeding, we create additional notations. We let $X$ be partitioned as $\left(X^{C},X^{D}\right)$, where $X^{C}$ and $X^{D}$ consist of covariates that are continuously and discretely distributed, respectively. Without loss of generality, we suppose that the vector of zeros is in the support of $X^{D}$. Denote $X_{k}$, $X_{k}^{C}$, and $X_{k}^{D}$ as the $k$-th coordinates of $X$, $X^{C}$, and $X^{D}$, respectively. Denote $x^{C}$ as a generic element in the support of $X^{C}$; likewise for $x^{D}$, $x_{k}^{C}$, and $x_{k}^{D}$. Moreover, denote $\pi_{0}\left(x^{C}\right)=\pi\left(x^{C},0\right)$, where the discrete covariates are equal to zero. If $\pi_{0}$ is differentiable at $x^{C}$, then we denote $\partial_{k}\pi_{0}\left(x^{C}\right)=\left.\partial\pi_{0}\left(x^{C}\right)\right/\partial x_{k}^{C}$ as its partial derivative with respect to the $k$-th argument. For any $x^{D}$, we denote $x^{Dk}$ as the $\dim\left(x^{D}\right)\times 1$ vector with the $k$-th coordinate being equal to $x_{k}^{D}$ and all the other coordinates being equal to zero.
Assumption L (Linearity). Assume that $E\left[\left.Y_{d}\right|X=x\right]=\alpha_{d}+x^{\prime}\beta_{d}$ for some fixed $\alpha_{d}$ and $\beta_{d}$, $d=0,1$.
Assumption NL (Non-Linearity). Assume that $\pi_{0}$ satisfies the following NL1 when $\dim\left(X^{C}\right)=1$ or NL2 when $\dim\left(X^{C}\right)\geq 2.$
– NL1 ($\dim\left(X^{C}\right)=1$): there exist two different constants $x^{C}$, $\tilde{x}^{C}$ in the support of $X^{C}$ such that $\pi_{0}\left(x^{C}\right)=\pi_{0}\left(\tilde{x}^{C}\right)$.
– NL2 ($\dim\left(X^{C}\right)\geq 2$): there exist two vectors $x^{C}$, $\tilde{x}^{C}$ in the support of $X^{C}$ and two elements $k$, $j$ in set $\left\{1,2,\cdots,\dim\left(X^{C}\right)\right\}$ such that $\pi_{0}$ and $E\left[Y\left|X^{C}=x^{C},X^{D}=0,D=d\right.\right]$, $d=0,1$, are differentiable at $x^{C}$ and $\tilde{x}^{C}$, and that (i) $\partial_{k}\pi_{0}\left(x^{C}\right)\neq 0$, (ii) $\partial_{j}\pi_{0}\left(x^{C}\right)\neq 0$, (iii) $\partial_{k}\pi_{0}\left(\tilde{x}^{C}\right)\neq 0$, (iv) $\partial_{j}\pi_{0}\left(\tilde{x}^{C}\right)\neq 0$, and (v) $\left.\partial_{k}\pi_{0}\left(x^{C}\right)\right/\partial_{j}\pi_{0}\left(x^{C}\right)\neq\left.\partial_{k}\pi_{0}\left(\tilde{x}^{C}\right)\right/\partial_{j}\pi_{0}\left(\tilde{x}^{C}\right)$.
@@
Assumption CMI (Conditional Mean Independence). Denote $U_{d}=Y_{d}-X^{\prime}\beta_{d}$, $d=0,1$. Assume that $E\left[U_{d}\left|V,X\right.\right]=E\left[U_{d}\left|V\right.\right]$ with probability one for $d=0,1$.
Assumption L imposes a linear restriction on the potential outcome regression functions, which is a common practice in empirical studies. The linear restriction may seem too strong compared with that in the existing identification strategies, which allow highly flexible (especially nonparametric) specifications on the potential outcomes. However, for the estimation of the MTE, the linear specification has been nearly universally adopted for tractability and interpretability (e.g., Kirkeboen et al., 2016; Kline and Walters, 2016; Brinch et al., 2017; Heckman et al., 2018; Mogstad et al., 2021; Aryal et al., 2022; Mountjoy, 2022). Noting that Assumption L implicitly requires the potential outcomes to be continuously distributed and supported on the entire real line, we would like to point out that our identification strategy can also be adapted to the case of limited valued outcomes by specifying linear latent index. A detailed discussion is left to Appendix A.3.
By contrast, Assumption NL requires the propensity score function to be nonlinear in the continuous covariates in a generalized sense. The combination of Assumptions L and NL will enable identification based on functional form in a semiparametric version that can realize the identification of linear coefficients by exploiting nonlinearity elsewhere in the model. The nonlinearity assumption has two different forms, depending on the number of continuous covariates. When two or more continuous covariates are available, Assumption NL2 will require some variation in $\pi_{0}$ to distinguish it from a linear index function. Concretely, NL2 will not hold if $\pi_{0}\left(x^{C}\right)=f\left(x^{C\prime}\gamma\right)$ for a smooth function $f$, because in this case, both sides of the inequality in (v) are equal to $\left.\gamma_{k}\right/\gamma_{j}$. Otherwise, however, it is difficult to construct examples that violate (v). In practice, (v) can be fulfilled even when $\pi_{0}$ is single-index specified, if interaction or/and quadratic terms are added, as shown in Example 3.
Example 3.
Consider the case of two continuous covariates. Suppose that for a smooth function $f$, $\pi_{0}\left(x^{C}\right)=f\left(\gamma_{1}x_{1}^{C}+\gamma_{2}x_{2}^{C}+\gamma_{3}x_{1}^{C}x_{2}^{C}\right)$ or $\pi_{0}\left(x^{C}\right)=f\left(\gamma_{1}x_{1}^{C}+\gamma_{2}x_{2}^{C}+\gamma_{3}\left(x_{1}^{C}\right)^{2}\right)$. Then, we obtain $\left.\partial_{1}\pi_{0}\left(x^{C}\right)\right/\partial_{2}\pi_{0}\left(x^{C}\right)=\left.\left(\gamma_{1}+\gamma_{3}x_{2}^{C}\right)\right/\left(\gamma_{2}+\gamma_{3}x_{1}^{C}\right)$ for the interaction case, or $\left.\partial_{1}\pi_{0}\left(x^{C}\right)\right/\partial_{2}\pi_{0}\left(x^{C}\right)=\left.\left(\gamma_{1}+2\gamma_{3}x_{1}^{C}\right)\right/\gamma_{2}$ for the quadratic case. In both cases, Assumption NL2 (v) will generally hold for $x^{C}$ and $\tilde{x}^{C}$ satisfying $x_{1}^{C}\neq\tilde{x}_{1}^{C}$.
Assumption NL2 requires the existence of at least two continuous covariates. However, in empirical studies based on survey data, most of the demographic characteristics are documented as discrete or categorical variables, such as age, gender, race, marital status, educational attainment, and so on. Therefore, we also impose Assumption NL1, as a supplement to NL2, to take into account the situation in which only one continuous covariate is available. NL1 requires the univariate function $\pi_{0}$ to be not one-to-one, but imposes no other smoothness or continuity restriction on $\pi_{0}$. This condition will hold if the probability of receiving treatment is unaffected by some change in the continuous covariate. A similar local irrelevance assumption is imposed in nonseparable models to attain point identification (Torgovitsky, 2015) or to test and relax the exclusion restriction (D’Haultfœuille et al., 2021). Following the line of Assumption NL1, our identification strategy works even when no continuous covariate is available in the data, and we put a discretely distributed covariate in $X^{C}$. In the extreme case of $X^{C}$ containing only one binary covariate, NL1 will be equivalent to the full irrelevance of $X^{C}$ to the treatment probability, which is a condition suggested by Chamberlain (1986) in the identification of the sample selection model. However, using only discrete covariates provides little identifying variation, which may lead to poor performance in the subsequent model estimation (Garlick and Hyman, 2022). Hence, we focus on the case of at least one continuous covariate. We also note that if $\pi_{0}$ is assumed to be differentiable as in NL2, then the local irrelevance condition can be rephrased as a simpler form and extended to multiple continuous covariates. Further discussion on NL1 and its variants is left to Appendix A.2.
The nonlinearity assumption does not exclude the widely specified linear-index treatment equation
$D=1\left\{X^{\prime}\gamma\geq U\right\},$
as long as the structural error $U$ is not independent of $X$. For instance, consider a multiplicatively heteroscedastic $U$ such that $U=\sigma\left(X\right)\tilde{U}$, where $\sigma\left(X\right)$ is a positive function and $\tilde{U}$ is an idiosyncratic error independent of $X$. The normalization derives
that is, the reduced-form error $V$ in (3) is equal to $\tilde{U}$, and the propensity score function is equal to $\pi\left(x\right)=\left.\left(x^{\prime}\gamma\right)\right/\sigma\left(x\right)$. In general, $\pi\left(x\right)$ is a nonlinear function. Another example is a linear model with endogeneity in a certain component $X^{e}$ of $X^{C}$, in which $\pi\left(x\right)=F_{U\left|X^{e}\right.}\left(\left.x^{\prime}\gamma\right|x^{e}\right)$. Given that $\pi\left(x\right)$ is generally highly nonlinear in $x^{e}$, Assumption NL1 holds straightforwardly, and NL2 holds with $x_{k}^{C}=x^{e}$. Unlike in the binary response model, the ubiquitous heteroscedasticity and endogeneity benefit our results while inducing no trouble, because the identification of the MTE is irrelevant to the structural coefficient $\gamma$. Our MTE is defined by the reduced-form treatment error; thus, all we need from the treatment equation is the propensity score, which has a reduced-form nature.
@@
Assumption CMI is standard in the MTE literature and commonly referred to as the separability or additive separability assumption (e.g., Brinch et al., 2017; Mogstad et al., 2018; Zhou and Xie, 2019), because it renders the MTE additively separable into observables and unobservables:
$\Delta^{\text{MTE}}\left(x,v\right)=x^{\prime}\left(\beta_{1}-\beta_{0}\right)+E\left[U_{1}-U_{0}\left|V=v\right.\right].$
@@
Therefore, this assumption implies that the shape of the MTE curve will not vary with covariates. By definition, $E\left[U_{d}\left|X\right.\right]=\alpha_{d}=E\left[U_{d}\right]$, and $V\perp\!\!\!\!\perp X$; hence, Assumption CMI essentially requires the conditional covariance of $U_{d}$ and $V$ to be independent of $X$, which is also a key assumption in the heteroscedasticity-based identification method (Lewbel, 2012). This assumption can provide an alternative source of exogenous excluded variation for the identification in place of the IV. It is implied by and much weaker than the full independence $\left(U_{d},U\right)\perp\!\!\!\!\perp X$, which is frequently imposed (often implicitly) in applied work, where $U$ is the structural treatment error in (1). In particular, Assumption CMI does not rule out the marginal dependence of $U_{d}$ or $U$ on $X$, as illustrated in Example 4. Assumption S imposes a mild support condition for technical reasons. A sufficient condition for Assumption S to hold is if $\pi_{0}\left(X^{C}\right)$ has full support on the unit interval.
@@
In this setting, $U$ is correlated to $X$. Through a property of the multivariate normal distribution, we obtain
@@
Under Assumptions L and CMI, we have
$E\left[Y\left|X=x,D=d\right.\right]=x^{\prime}\beta_{d}+g_{d}\left(\pi\left(x\right)\right)$ (5)
for $d=0,1$, where
@@
Several calculations show that
$E\left[\left.U_{0}\right|V=p\right]$ $=$ $g_{0}\left(p\right)-\left(1-p\right)g_{0}^{\left(1\right)}\left(p\right),$
$E\left[\left.U_{1}\right|V=p\right]$ $=$ $g_{1}\left(p\right)+pg_{1}^{\left(1\right)}\left(p\right),$
where $g_{d}^{\left(1\right)}$ denotes the derivative function of $g_{d}$. Therefore, the identification of $\Delta^{\text{MTE}}\left(x,v\right)$ can be obtained by identifying $\beta_{d}$ and $g_{d}\left(p\right)$ separately for $d=0,1$. The theorem below establishes this identification result via a construction method.
@@
If Assumptions L, NL, CMI, and S hold, then $\beta_{d}$ and $g_{d}\left(p\right)$ at all $p$ in the support of the propensity score $P$ are identified for $d=0,1$.
The proof of Theorem 1 is grounded on regression function (5), which summarizes the information from the data. As $g_{d}$ is unknown, we need to eliminate it through some subtraction to realize the identification of $\beta_{d}$. When only one continuous covariate exists, the subtraction will be straightforward according to Assumption NL1. Otherwise, our strategy is to perturb two continuous covariates, such as $x_{k}^{C}$ and $x_{j}^{C}$, in such a way that $\pi\left(x\right)$ remains unchanged. Specifically, we increase $x_{k}^{C}$ by a small $\epsilon$ and simultaneously change $x_{j}^{C}$ by $\epsilon$ multiplied by $\left.-\partial_{k}\pi_{0}\left(x^{C}\right)\right/\partial_{j}\pi_{0}\left(x^{C}\right)$, resulting in a perturbed value of the regression function. Subtracting the perturbed regression function from the original (5) will cancel out $g_{d}$ due to the equality of $\pi\left(x\right)$, giving rise to an equation for $\beta_{d,k}^{C}$ and $\beta_{d,j}^{C}$. Note that the multiplier $\left.-\partial_{k}\pi_{0}\left(x^{C}\right)\right/\partial_{j}\pi_{0}\left(x^{C}\right)$ is the partial derivative of $x_{j}^{C}$ with respect to $x_{k}^{C}$ if we view $x_{j}^{C}$ as an implicit function of the other continuous covariates by equating $\pi_{0}\left(x^{C}\right)$ to a constant. Accordingly, Assumption NL2 (v) implies two linearly independent equations, ensuring an exact solution (i.e., identification) of $\beta_{d,k}^{C}$ and $\beta_{d,j}^{C}$. The detailed proof of Theorem 1 is presented in Appendix A.1. It is worth mentioning that our strategy naturally features overidentification in the sense that generally more than two pairs of points in the support of $X^{C}$ satisfy Assumption NL, because $X^{C}$ is continuously distributed. As a result, the identification can also be represented as the average of solutions over all pairs of points satisfying Assumption NL.
Theorem 1 implies the identification of the MTE in the absence of IVs. Specifically,
This result allows practitioners to include all the relevant observed characteristics into both treatment and outcome equations, without imposing any exclusion or full independence assumptions. Under Theorem 1, the conventional causal parameters can also be identified without an instrument, provided that the support of $P$ contains 0 and/or 1 (which implies the identifiability of $g_{d}\left(0\right)$ and/or $g_{d}\left(1\right)$):

### v2

@@ 3 Identification
Our identification strategy relies on a linearity restriction on the potential outcome equations and a nonlinearity restriction on the propensity score function. The intuition is that the propensity score minus the linear outcome index will provide the excluded variation that perturbs treatment status, which plays the role of a continuous IV. Furthermore, to ensure the exogeneity of the excluded variation, it is necessary to impose a certain form of independence between the potential outcome residuals and covariates. We denote $h_{d}\left(x\right)=E\left[\left.Y_{d}\right|X=x\right]$ and $U_{d}=Y_{d}-h_{d}\left(X\right)$, $d=0,1$, as the regression functions and residuals of potential outcomes.
@@
Assumption CMI (Conditional Mean Independence). Assume that $E\left[U_{1}-U_{0}\left|V,X\right.\right]=E\left[U_{1}-U_{0}\left|V\right.\right]$ with probability one.
@@
Assumption CMI is standard in the MTE literature and commonly referred to as separability or additive separability assumption (e.g., Brinch et al., 2017; Mogstad et al., 2018; Zhou and Xie, 2019), because it’s a necessary and sufficient condition for the MTE to be additively separable into observables and unobservables (Zhou and Xie, 2019, p.3074):
$\Delta^{\text{MTE}}\left(x,v\right)$ $=$ $h_{1}\left(x\right)-h_{0}\left(x\right)+E\left[U_{1}-U_{0}\left|X=x,V=v\right.\right]$
$=$ $h_{1}\left(x\right)-h_{0}\left(x\right)+E\left[U_{1}-U_{0}\left|V=v\right.\right].$
@@
Namely, under Assumption CMI, the shape of the MTE curve with respect to $v$ will not vary with covariates. Assumption CMI is partly justified by $E\left[U_{d}\left|X\right.\right]=0$ and $V\perp\!\!\!\!\perp X$, which come directly from the definition. On this basis, it is sufficient for Assumption CMI to hold if the conditional covariance of $U_{d}$ and $V$ is independent of $X$, which is a key assumption in the heteroscedasticity-based identification method as well (Lewbel, 2012). Assumption CMI is implied by and much weaker than the full independence $\left(U_{d},U\right)\perp\!\!\!\!\perp X$ frequently imposed (often implicitly) in applied work, where $U$ is the structural treatment error in (1). In particular, Assumption CMI doesn’t rule out the marginal dependence of $U_{d}$ or $U$ on $X$, as illustrated by Example 3.
@@
In this setting, $U$ is correlated to $X$ with an unconstrained correlation coefficient. Through a property of the multivariate normal distribution, we obtain
@@
For the purpose of identifying the MTE, we first establish the relationship between $\Delta^{\text{MTE}}\left(x,v\right)$ and observed regression functions. Under Assumption CMI, the observed regression functions for treated and untreated individuals are additively separable in a similar way:
$E\left[Y\left|X,D=d\right.\right]$ $=$ $E\left[Y_{d}\left|X,D=d\right.\right]$ (5)
$=$ $h_{d}\left(X\right)+E\left[U_{d}\left|X,D=d\right.\right]$
$=$ $h_{d}\left(X\right)+g_{d}\left(P\right),\text{ }d=0,1,$
where
@@
By multiplying both sides of the above equations with $p$ or $1-p$ in order to eliminate the denominators and then differentiating with respect to $p$, we obtain
$g_{1}\left(p\right)+pg_{1}^{\left(1\right)}\left(p\right)$ $=$ $E\left[\left.U_{1}\right|V=p\right],$
$-g_{0}\left(p\right)+\left(1-p\right)g_{0}^{\left(1\right)}\left(p\right)$ $=$ $-E\left[\left.U_{0}\right|V=p\right],$
where $g_{d}^{\left(1\right)}$ denotes the derivative function of $g_{d}$. Therefore, the MTE can be represented as
$\Delta^{\text{MTE}}\left(x,v\right)=h_{1}\left(x\right)-h_{0}\left(x\right)+\left[g_{1}\left(v\right)-g_{0}\left(v\right)\right]+vg_{1}^{\left(1\right)}\left(v\right)+\left(1-v\right)g_{0}^{\left(1\right)}\left(v\right),$ (6)
and the identification of it can be achieved by identifying $h_{d}$ and $g_{d}$ from the regression equation (5).
Recall that $P=\pi\left(X\right)$, where $\pi$ is a deterministic function, albeit unknown. Given that the argument of $g_{d}$ in (5) exhibits no variation apart from that generated by $X$, it is impossible to distinguish between $g_{d}$ and $h_{d}$ in the absence of additional assumptions. For example, one may choose to let $h_{d}$ absorb $g_{d}\left(\pi\left(\cdot\right)\right)$ and set $g_{d}=0$, or vice versa. Functional form restrictions that differentiate $h_{d}$ and $g_{d}\left(\pi\left(\cdot\right)\right)$ from one another can address this issue and facilitate the identification of both $h_{d}$ and $g_{d}$. To this end, we impose linearity on $h_{d}$ and nonlinearity on $\pi$.
Assumption L (Linearity). Assume that $E\left[\left.Y_{d}\right|X\right]=\alpha_{d}+X^{\prime}\beta_{d}$ with probability one for some fixed $\alpha_{d}$ and $\beta_{d}$, $d=0,1$.
Assumption L imposes a linear restriction on the potential outcome regression functions, which is a common practice in empirical studies. The linear restriction may seem too strong compared with that in the existing identification strategies, which allow highly flexible (especially nonparametric) specifications on the potential outcomes. However, for the estimation of MTE, the linear specification has been nearly universally adopted for tractability and interpretability (e.g., Kirkeboen et al., 2016; Kline and Walters, 2016; Brinch et al., 2017; Heckman et al., 2018; Mogstad et al., 2021; Aryal et al., 2022; Mountjoy, 2022). Noting that Assumption L implicitly requires the potential outcomes to be continuously distributed and supported on the entire real line, we would like to point out that our identification strategy can also be adapted to the case of limited valued outcomes by specifying linear latent index. A detailed discussion is left to Appendix C. With a little abuse of notation, we redenote $U_{d}$ to absorb the intercept $\alpha_{d}$ so that
$Y_{d}=X^{\prime}\beta_{d}+U_{d},$
and thus that the regression equation (5) and the MTE become
$E\left[Y\left|X,D=d\right.\right]$ $=$ $X^{\prime}\beta_{d}+g_{d}\left(P\right),\text{ }d=0,1,$ (7)
To present the nonlinearity assumption, we let $X$ be partitioned as $\left(X^{C},X^{D}\right)$, where $X^{C}$ and $X^{D}$ consist of covariates that are continuously and discretely distributed, respectively. We denote $X_{k}$, $X_{k}^{C}$, and $X_{k}^{D}$ as the $k$-th coordinates of $X$, $X^{C}$, and $X^{D}$, respectively. And we denote $x^{C}$ as a generic element in the support of $X^{C}$; likewise for $x^{D}$, $x_{k}^{C}$, and $x_{k}^{D}$. The nonlinearity assumption requires the propensity score function $\pi$ to be nonlinear in $X^{C}$ given a benchmark value of $X^{D}$. Without loss of generality, we suppose that the vector of zeros is in the support of $X^{D}$ and is the benchmark value. We denote $\pi_{0}\left(x^{C}\right)=\pi\left(x^{C},0\right)$ for notational convenience, where the discrete covariates are equal to zero.
Assumption NL (Non-Linearity). Assume that $\pi_{0}\left(x^{C}\right)$ and $E\left[Y\left|X^{C}=x^{C},X^{D}=0,D=d\right.\right]$, $d=0,1$, are differentiable with respect to $x^{C}$, and that the derivative of $\pi_{0}$ satisfies the following NL2 when $\dim\left(X^{C}\right)\geq 2$ or NL1 when $\dim\left(X^{C}\right)=1.$
– NL2 ($\dim\left(X^{C}\right)\geq 2$): there exist two vectors $x^{C}$, $\tilde{x}^{C}$ in the support of $X^{C}$ and two elements $k$, $j$ in set $\left\{1,2,\cdots,\dim\left(X^{C}\right)\right\}$ such that (i) $\partial_{k}\pi_{0}\left(x^{C}\right)\neq 0$, (ii) $\partial_{j}\pi_{0}\left(x^{C}\right)\neq 0$, (iii) $\partial_{k}\pi_{0}\left(\tilde{x}^{C}\right)\neq 0$, (iv) $\partial_{j}\pi_{0}\left(\tilde{x}^{C}\right)\neq 0$, and (v) $\left.\partial_{k}\pi_{0}\left(x^{C}\right)\right/\partial_{j}\pi_{0}\left(x^{C}\right)\neq\left.\partial_{k}\pi_{0}\left(\tilde{x}^{C}\right)\right/\partial_{j}\pi_{0}\left(\tilde{x}^{C}\right)$, where $\partial_{k}\pi_{0}\left(x^{C}\right)=\left.\partial\pi_{0}\left(x^{C}\right)\right/\partial x_{k}^{C}$ is the partial derivative of $\pi_{0}$ with respect to the $k$-th argument.
– NL1 ($\dim\left(X^{C}\right)=1$): there exists a constant $\tilde{x}^{C}$ in the support of $X^{C}$ such that $\pi_{0}^{\left(1\right)}\left(\tilde{x}^{C}\right)=0$, where $\pi_{0}^{\left(1\right)}\left(x^{C}\right)=\left.d\pi_{0}\left(x^{C}\right)\right/dx^{C}$ is the univariate derivative of $\pi_{0}$.
Assumption NL requires the propensity score function to be nonlinear in continuous covariates in a generalized sense. The combination of Assumptions L and NL will enable identification based on functional form in a semiparametric version, which can realize the identification of linear coefficients by exploiting nonlinearity elsewhere in the model. The nonlinearity assumption has two different forms, depending on the number of continuous covariates. When two or more continuous covariates are available, Assumption NL2 will require some variation in $\pi_{0}$ to distinguish it from a linear-index function. Concretely, NL2 will not hold if $\pi_{0}\left(x^{C}\right)=f\left(x^{C\prime}\gamma\right)$ for a smooth function $f$, because in this case, both sides of the inequality in (v) are equal to $\left.\gamma_{k}\right/\gamma_{j}$. Otherwise, however, it is difficult to construct examples that violate (v). In practice, (v) can be fulfilled even when $\pi_{0}$ is single-index specified, if interaction or/and quadratic terms are added, as shown in Example 4.
Example 4.
Consider the case of two continuous covariates. Suppose that for a smooth function $f$, $\pi_{0}\left(x^{C}\right)=f\left(\gamma_{1}x_{1}^{C}+\gamma_{2}x_{2}^{C}+\gamma_{3}x_{1}^{C}x_{2}^{C}\right)$ or $\pi_{0}\left(x^{C}\right)=f\left(\gamma_{1}x_{1}^{C}+\gamma_{2}x_{2}^{C}+\gamma_{3}\left(x_{1}^{C}\right)^{2}\right)$. Then, we obtain $\left.\partial_{1}\pi_{0}\left(x^{C}\right)\right/\partial_{2}\pi_{0}\left(x^{C}\right)=\left.\left(\gamma_{1}+\gamma_{3}x_{2}^{C}\right)\right/\left(\gamma_{2}+\gamma_{3}x_{1}^{C}\right)$ for the interaction case, or $\left.\partial_{1}\pi_{0}\left(x^{C}\right)\right/\partial_{2}\pi_{0}\left(x^{C}\right)=\left.\left(\gamma_{1}+2\gamma_{3}x_{1}^{C}\right)\right/\gamma_{2}$ for the quadratic case. In both cases, Assumption NL2.(v) will generally hold for $x^{C}$ and $\tilde{x}^{C}$ satisfying $x_{1}^{C}\neq\tilde{x}_{1}^{C}$.
Assumption NL2 requires the existence of at least two continuous covariates. However, in empirical studies based on survey data, most of the demographic characteristics are documented as discrete or categorical variables, such as age, gender, race, marital status, educational attainment, and so on. Therefore, we also impose Assumption NL1, as a supplement to NL2, to take into account the situation in which only one continuous covariate is available. NL1 requires $\pi_{0}$ to have a stationary point, which implies that $\pi_{0}$ is necessarily nonlinear. NL1 will hold if the probability of receiving treatment is unaffected by some local change in the continuous covariate. It’s a differential version of the local irrelevance assumption imposed in nonseparable models for attaining point identification (e.g., Torgovitsky, 2015; D’Haultfœuille et al., 2021). NL1 can be extended to the case of no continuous covariate, as discussed in Appendix B. However, using only discrete covariates provides little identifying variation, which may lead to poor performance in the subsequent model estimation (Garlick and Hyman, 2022). Hence, we focus on the case of at least one continuous covariate.
We note that Assumption NL doesn’t exclude the widely specified linear-index treatment equation, as long as the structural error is not independent of covariates.
Example 5.
Suppose the treatment status is determined by a linear-index threshold crossing rule:
$D=1\left\{X^{\prime}\gamma\geq U\right\},$
where $U$ is not independent of $X$. Namely, suppose $\mu\left(X\right)=X^{\prime}\gamma$ in the structural treatment equation (1). Consider a multiplicatively heteroscedastic $U$ such that $U=\sigma\left(X\right)\tilde{U}$, where $\sigma\left(X\right)$ is a positive function and $\tilde{U}$ is an idiosyncratic error independent of $X$. The normalization derives
that is, the reduced-form error $V$ in (3) is equal to $\tilde{U}$, and the propensity score function is equal to $\pi\left(x\right)=\left.\left(x^{\prime}\gamma\right)\right/\sigma\left(x\right)$. In general, $\pi\left(x\right)$ is a nonlinear function. Another specification we consider is a linear-index model with endogeneity in a certain component $X^{e}$ of $X^{C}$, in which $\pi\left(x\right)=F_{U\left|X^{e}\right.}\left(\left.x^{\prime}\gamma\right|x^{e}\right)$. Given that $\pi\left(x\right)$ is generally highly nonlinear in $x^{e}$, Assumption NL1 holds straightforwardly, and NL2 holds with $x_{k}^{C}=x^{e}$.
Unlike in the binary response model, the ubiquitous heteroscedasticity and endogeneity benefit our results while inducing no trouble, because the identification of the MTE is irrelevant to the structural coefficients in $\gamma$. Our MTE is defined by the reduced-form treatment error; thus, all we need from the treatment equation is the propensity score, which has a reduced-form nature.
Assumption NL ensures identification of coefficients of continuous covariates. In order to identify coefficients of discrete covariates, we impose a mild support condition. For any $x_{k}^{D}\neq 0$, we denote $x^{Dk}$ as the $\dim\left(X^{D}\right)\times 1$ vector with the $k$-th coordinate being equal to $x_{k}^{D}$ and all the other coordinates being equal to zero.
A sufficient condition for Assumption S to hold is that $\pi_{0}\left(X^{C}\right)$ has a full support on the unit interval, or, more generally, that the support of $\pi_{0}\left(X^{C}\right)$ is overlapped with those of $\pi\left(X^{C},x^{D}\right)$ for all $x^{D}\neq 0$. Otherwise, we have to find an $x_{k}^{D}\neq 0$ for each $k$ such that the support of $\pi\left(X^{C},x^{Dk}\right)$ is overlapped with that of $\pi_{0}\left(X^{C}\right)$. The following theorem establishes our main identification result.
@@
If Assumptions CMI, L, NL, and S hold, then $\beta_{d}$ and $g_{d}\left(p\right)$ at all $p$ in the support of the propensity score $P$ are identified for $d=0,1$.
The proof of Theorem 1 is grounded on the observed regression functions (7), which summarize the information from the data. As $g_{d}$ is unknown, we need to eliminate it through some subtraction to realize the identification of $\beta_{d}$. When only one continuous covariate exists, the subtraction can be carried out locally around the point satisfying Assumption NL1. Otherwise, our strategy is to perturb two continuous covariates, such as $x_{k}^{C}$ and $x_{j}^{C}$, in such a way that $\pi\left(x\right)$ remains unchanged. Specifically, for each group $d$, we increase $x_{k}^{C}$ by a small $\epsilon$ and simultaneously change $x_{j}^{C}$ by $\epsilon$ multiplied by $\left.-\partial_{k}\pi_{0}\left(x^{C}\right)\right/\partial_{j}\pi_{0}\left(x^{C}\right)$, resulting in a perturbed value of the regression function. Subtracting the perturbed regression function from the original (7) will cancel out $g_{d}$ due to the equality of $\pi\left(x\right)$, giving rise to an equation for $\beta_{d,k}^{C}$ and $\beta_{d,j}^{C}$. Note that the multiplier $\left.-\partial_{k}\pi_{0}\left(x^{C}\right)\right/\partial_{j}\pi_{0}\left(x^{C}\right)$ is the partial derivative of $x_{j}^{C}$ with respect to $x_{k}^{C}$ if we view $x_{j}^{C}$ as an implicit function of the other continuous covariates by equating $\pi_{0}\left(x^{C}\right)$ to a constant. Accordingly, Assumption NL2.(v) implies two linearly independent equations, ensuring an exact solution (i.e., identification) of $\beta_{d,k}^{C}$ and $\beta_{d,j}^{C}$. The detailed proof of Theorem 1 is presented in Appendix A. It is worth mentioning that our strategy naturally features overidentification in the sense that generally more than one pair of points in the support of $X^{C}$ satisfy Assumption NL2 or more than one point satisfies NL1, because $X^{C}$ is continuously distributed. Moreover, there may be more than one value of $X^{D}$ that satisfies Assumptions NL and S. Consequently, the identification can also be represented as the average of solutions over all points of $X^{C}$ satisfying Assumption NL and over all values of $X^{D}$ satisfying Assumptions NL and S.
According to (8), Theorem 1 implies the identification of MTE in the absence of IVs. Specifically, This result allows practitioners to include all the relevant observed characteristics into both treatment and outcome equations, without imposing any exclusion or full independence assumptions. Under Theorem 1, the conventional causal parameters can also be identified without instruments, provided that the support of $P$ contains 0 and/or 1 (which implies identifiability of $g_{d}\left(0\right)$ and/or $g_{d}\left(1\right)$):